OpenStudy (anonymous):

square root of 28- the square root of 112 +3 square root 63

OpenStudy (lgbasallote):

hint: \[\sqrt 28 \implies \sqrt{4 \times 7} \implies 2 \sqrt 7\] \[\sqrt{112} \implies \sqrt{16 \times 7} \implies 4\sqrt 7\] \[\sqrt{63} \implies \sqrt{9 \times 7} \implies 3\sqrt 7\] does that help?

OpenStudy (anonymous):

yea... somehow i got 7 root 7

OpenStudy (lgbasallote):

nope. what is the value of \[2\sqrt 7 - 4\sqrt 7 + 3\sqrt 7\]

OpenStudy (anonymous):

is it just root 7

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

ok...but does that include the 3 before the root 63?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (lgbasallote):

oh wait.. i did not see that o.O

OpenStudy (lgbasallote):

sorry \(7\sqrt 7\) is right then

OpenStudy (anonymous):

the problem is root 28 - root 112 + 3 root 63

OpenStudy (anonymous):

are you serious?

OpenStudy (anonymous):

well thanks for the explanation, it helped!!!!